# HiSAC: High-Resolution Sensing with Multiband Communication Signals

### v1

@@ 7. Related Work
Super-resolution wireless sensing. Subspace-based super-resolution methods (Kung et al., 1983) and CS (Eldar and Kutyniok, 2012) are widely used in radar and ISAC, but are still limited by the transmission bandwidth. Some of their recent applications to ISAC are found in (Liu et al., 2020; Zhang et al., 2023). Other approaches have employed the spatial diversity of the RX array (Zhang et al., 2021b) to perform high-resolution imaging in the mmWave band, but they can not be applied to low-cost systems with a single RF chain such as analog beamforming systems. In (Tagliaferri et al., 2024; Manzoni et al., 2023), a novel approach to apply distributed Synthetic Aperture Radar (SAR) techniques to vehicular sensing systems is introduced. This approach boosts the resolution by combining distributed devices, thus requiring multiple cooperating nodes. HiSAC instead reuses the frequency diversity of communication systems without additional requirements. Notably, HiSAC could be combined with any of the above techniques to enhance their resolution.
@@
Multiband radar sensing. Multiband sensing has been studied in the radar literature. (Cuomo et al., 1999) has first proposed bandwidth interpolation between two subbands to increase ranging resolution via Auto-Regressive (AR) modeling and non-linear optimization. Other works have followed a similar research direction adopting different algorithms for combining the subbands (Zhang and Chen, 2014; Hussain et al., 2018; Xiong et al., 2017; Jia and Xu, 2006; Tian et al., 2013; Zhang et al., 2017; van Dorp et al., 2010). The above works are based on radar systems, which adopt optimized chirp waveforms and can rely on wide individual subbands. This significantly simplifies the problem with respect to an ISAC setting where CFR estimates are not under control and can be very narrow, which makes radar approaches underperform. To solve this problem, HiSAC innovates with a progressive combination of subbands over coherent subsystems first, and then over the full band of interest.
@@
OFDM-based multiband processing. Several works have demonstrated that combining multiple frequency bands can boost the resolution of OFDM systems in active localization. In (Kotaru et al., 2015; Xie et al., 2015; Xiong et al., 2015), SpotFi, Splicer, and ToneTrack are presented, which combine (stitch) multiple contiguous or overlapping Wi-Fi subbands to increase the multipath resolution. To eliminate phase offsets, they use linear fitting of the unwrapped phase, which is prone to errors when used on non-contiguous, narrow subbands spanning several GHz. Other Wi-Fi-based approaches (Khalilsarai et al., 2019, 2020; Vasisht et al., 2016) have tackled the same problem but rely on a handshaking process between TX and RX to eliminate phase offsets, which does not apply to passive sensing in ISAC. (Noschese et al., 2020; Wan et al., 2022, 2023; Kazaz et al., 2021) propose alternative algorithms based on maximum-likelihood. All these approaches target active localization, in which the RX device is the localized target. HiSAC instead localizes targets from backscattered reflections, as done by radar systems.
@@
More recently, (Wan et al., 2024) has brought the research attention to exploiting multiband CFR to perform wideband radar-like ranging in ISAC. To the best of our knowledge, the only system that tackles this problem is UWB-Fi (Wang et al., 2024), in the sub-6 GHz unlicensed spectrum. However, this system is based on a neural network that learns to combine subbands and compensate for phase offsets. Changing hardware, frequency band, or technology (OFDM vs SC) may require retraining the system which is time-consuming and requires new data. Conversely, HiSAC does not require training and generalizes to different implementations.
@@
Cross-band channel prediction. A recent line of work has investigated channel prediction in one frequency band from an available channel estimate in a different band (Bakshi et al., 2019; Cho et al., 2023; Li et al., 2023; Liu et al., 2021; Banerjee et al., 2024). Although the general idea of this problem is linked to the multiband setting, predicting the channel in a different band is significantly different from combining multiple non-contiguous incoherent CFR estimates to increase the total system bandwidth. Moreover, the above works focus on communication, hence the channel prediction is mostly aimed at estimating SNR at the RX. Conversely, HiSAC is an ISAC system that estimates fine-grained complex amplitudes and delays of individual paths in the channel.
@@ 9. Conclusion
The problem of achieving range super-resolution in ISAC systems with narrow and discontinuous subbands is addressed in this paper. To solve it, we present HiSAC, a general signal-processing method for coherent multiband ranging that enhances the resolution of existing communication systems by only reusing channel estimates obtained via pilot signals. Our approach does not rely on a specific hardware or protocol and works across different communication technologies (i.e., 5G-NR or IEEE 802.11ay). Our extensive experiments with objects and humans demonstrate that HiSAC enhances the resolution by up to 20 times compared to single-band processing, occupying the same bandwidth in each time slot.

### v2

@@ 7. Related Work
Super-resolution wireless sensing. Subspace-based super-resolution methods (Kung et al., 1983) and compressed sensing (Eldar and Kutyniok, 2012) are widely used in radar and ISAC, but are still limited by the transmission bandwidth. Some of their recent applications to ISAC are found in (Liu et al., 2020; Zhang et al., 2023). Other approaches have employed the spatial diversity of the RX array (Zhang et al., 2021c) to perform high-resolution imaging in the mmWave band, but they can not be applied to low-cost systems with a single RF chain such as analog beamforming systems. In (Tagliaferri et al., 2024; Manzoni et al., 2023), a novel approach is introduced to apply distributed Synthetic Aperture Radar (SAR) techniques to vehicular sensing systems. This approach boosts the resolution by combining distributed devices, thus requiring multiple cooperating nodes. HiSAC instead reuses the frequency diversity of communication systems without additional requirements. Notably, HiSAC could be combined with any of the above techniques to enhance their resolution.
@@
Multiband radar sensing. Multiband sensing has been studied in the radar literature. (Cuomo et al., 1999) has first proposed bandwidth interpolation between two subbands to increase ranging resolution via Auto-Regressive (AR) modeling and non-linear optimization. Other works have followed a similar research direction adopting different algorithms for combining the subbands (Zhang and Chen, 2014; Hussain et al., 2018; Xiong et al., 2017; Jia and Xu, 2006; Tian et al., 2013; Zhang et al., 2017; van Dorp et al., 2010). The above works are based on radar systems with optimized chirp waveforms using wide individual subbands. This significantly simplifies the problem with respect to an ISAC setting where CFR estimates are not under control and can be very narrow, making radar approaches underperform. To solve this problem, HiSAC innovates with a progressive combination of subbands over coherent subsystems first, and then over the full band of interest.
@@
OFDM-based multiband processing. Several works have demonstrated that combining multiple frequency bands can boost the resolution of OFDM systems in active localization. In (Kotaru et al., 2015; Xie et al., 2015; Xiong et al., 2015), SpotFi, Splicer, and ToneTrack are presented, which combine (stitch) multiple contiguous or overlapping Wi-Fi subbands to increase the multipath resolution. They use linear fitting of the unwrapped phase to eliminate phase offsets, which leads to errors when used on non-contiguous, narrow subbands spanning several GHz. Other Wi-Fi-based approaches (Khalilsarai et al., 2019, 2020; Vasisht et al., 2016) have tackled the same problem but rely on a handshaking process between TX and RX to eliminate phase offsets, which does not apply to passive sensing in ISAC. Similarly, also the frequency hopping strategies used in (Bansal et al., 2021; Chen et al., 2021) require the cooperation of the receiver node, which may not be available in ISAC. (Noschese et al., 2020; Wan et al., 2022, 2023; Kazaz et al., 2021) propose alternative algorithms based on maximum-likelihood. All these approaches target active localization, in which the RX device is the localized target. HiSAC instead localizes targets from backscattered reflections, as done by radars.
@@
More recently, (Wan et al., 2024) has brought the research attention to exploiting multiband CFR to perform wideband radar-like ranging in ISAC. To the best of our knowledge, the only system that tackles this problem is UWB-Fi (Li et al., 2024), in the sub-6 GHz unlicensed spectrum. However, this system is based on a neural network that learns to combine subbands and compensate for phase offsets. Changing hardware, frequency band, or technology (OFDM vs SC) may require retraining the system which is time-consuming and requires new data. Conversely, HiSAC does not require training and generalizes to different implementations.
@@
Cross-band channel prediction. A recent line of work has investigated channel prediction in one frequency band from an available channel estimate in a different band (Bakshi et al., 2019; Cho et al., 2023; Li et al., 2023; Liu et al., 2021; Banerjee et al., 2024; Shepard et al., 2012). This is typically done to avoid having to transmit feedback channel information in downlink/uplink, thus reducing overhead. Although the general idea of this problem is linked to the multiband setting, predicting the channel in a different band is significantly different from combining multiple non-contiguous incoherent CFR estimates to increase the total system bandwidth. Moreover, the above works focus on communication, hence the channel prediction is mostly aimed at estimating SNR at the RX. Conversely, HiSAC is an ISAC system that estimates fine-grained complex amplitudes and delays of individual paths in the channel.
@@ 9. Conclusion
The problem of achieving range super-resolution in ISAC systems with narrow and discontinuous subbands is addressed in this paper. To solve it, we present HiSAC, a general signal-processing method for coherent multiband ranging that enhances the resolution of existing communication systems by only reusing channel estimates obtained via pilot signals. Our approach does not rely on specific hardware or protocol and works across different communication technologies (i.e., 5G-NR or IEEE 802.11ay). Our extensive experiments with objects and humans demonstrate that HiSAC enhances the resolution by up to 20 times compared to single-band processing, occupying the same bandwidth in each time slot.
Acknowledgements.
The work of Jacopo Pegoraro and Michele Rossi was funded by the European Union under the Italian National Recovery and Resilience Plan (NRRP) Mission 4, Component 2, Investment 1.3, CUP C93C22005250001, partnership on “Telecommunications of the Future” (PE00000001 - program “RESTART”). The work of Jesus O. Lacruz and Joerg Widmer has been partially funded by project PID2022-136769NB-I00 funded by MCIN/AEI /10.13039/501100011033 / FEDER, UE, and by project RISC-6G TSI-063000-2021-59, funded by Spanish Ministry of Economic Affairs and Digital Transformation, European Union NextGeneration-EU.